Marty leaves the airport in his private plane and flies due east at 186 mph. Two hours later, a jet leaves the same airport and flies due east at 434 mph. When will the jet overtake Marty's plane?
Answer by stanbon(75887)

Marty leaves the airport in his private plane and flies due east at 186 mph. Two hours later, a jet leaves the same airport and flies due east at 434 mph. When will the jet overtake Marty's plane?
-------------------------------
1st plane DATA:
rate = 186 mph ; time = x hrs ; distance = 186x miles
-------------------------------
2nd plane DATA:
rate = 434 mph; time = x-2 hrs; distance = 434(x-2) miles
-------------------------------
EQUATION:
distance = distance
186x = 434x - 2*434
-248x = -2*434
x = 3 1/2 hours (time the 1st plane is in the air)
x-2= 1 1/2 hours (time the 2nd plane takes to overtake the 1st plane)
===========================================
Cheers,
Stan H.